OpenStudy (cedric):

Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 11 3x + 4y = 24

OpenStudy (anonymous):

no

OpenStudy (cedric):

?

OpenStudy (anonymous):

Twice Bob's age + Alice's age = 11; Thrice Bob's age + 4x Alice's age = 24; find their ages?

OpenStudy (cedric):

ok how could i find their ages? bc 3 times bobs age= + 4 times alices age = 24 like idk how to find them

OpenStudy (anonymous):

To solve, rearrange the first equation: y=11-2x then substitute y into the second equation. 3x + 4(11-2x) = 24 3x+44-8x = 24 -5x = -20 x = 4 So Bob is age 4. sub in known value into either equation. 2(4)+y=11 8+y=11 y=3 So Alice is age 3. Then check: 2(4)+(3) is equal to 11 3(4)+4(3) is equal to 24. So there is the solution

OpenStudy (cedric):

oh ok i see now, thank you

OpenStudy (anonymous):

no probs ;)